OpenStudy (anonymous):

Solve each equation for y and identify the slope and y- intercept. Type any "mx" portion expressed as a fraction in the following manner. Example: y = 7x/8 + 4 x + y = 6 y = ____ + 6 m = b =

OpenStudy (the_fizicx99):

ok, simply subtract x from both sides in x + y = 6 Then you should be left with y = -x + 6 the slope is -1x, -1x is the same thing as -x... the y intercept in y = mx + b is B